If the weight of 5.6 lit of of a gas is 11g at stp.The gas is


1. SO2


2. NO


3. N2O

Grade:10

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Amit

wt of 22.4 lit gas=(22.4/5.6)*11=44g

wt of N2O = 14*2+16=44g

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

vikas askiitian expert
13 years ago

22.4L of gas = 1mol

1L= 1/22.4 mol

5.6L = 1/4 mol

let molar mass of gas is M then

5.6L = M/4gm =11gm  (given)

    so M=11*4=44gm

so the gas is N2O